OpenStudy (anonymous):

please help :) State the degree: rst+r^2s^2+r^3s^2

Mehek (mehek14):

options?

OpenStudy (anonymous):

it's free response :/

Mehek (mehek14):

rst has a degree of 3 because there's only variables so the exponent is 1 for each

Mehek (mehek14):

\(r^2s^2\) has 4 because add the exponents 2+2 = 4

Mehek (mehek14):

and the last one has the exponents 3 and 2 so 3+2 = 5

Mehek (mehek14):

I don't know if i'm supposed to combine the degrees

Mehek (mehek14):

are you supposed to put a number as the answer?

OpenStudy (anonymous):

yep!

Mehek (mehek14):

so add all of the degrees?

OpenStudy (anonymous):

would it be 10?

Mehek (mehek14):

3+4+5=12

OpenStudy (anonymous):

oh oops, thanks! :)

Mehek (mehek14):

yw \(\ddot\smile\)